find the equation of the ellipse with the following properties.
x-intercepts (2,0) and (-2,0); y-intercepts (0,4) and (0,-4)

Answer by lwsshak3(11628)

find the equation of the ellipse with the following properties.
x-intercepts (2,0) and (-2,0); y-intercepts (0,4) and (0,-4)
**
This is an ellipse with vertical major axis.
Its standard form of equation: (x-h)^2/b^2+(y^2)/a^2=1, (a>b), (h,k)=(x,y) coordinates of center
center: (0,0)
length of vertical major axis=8=2a
a=4
a^2+16
length of minor axis=4=2b
b=2
b^2=4
Equation:
x^2/4+y^2/16=1